OpenStudy (anonymous):

solving fractions. Solve for x. 4/3x + 2/3 = 6

OpenStudy (mkmkasim):

x should = 5

OpenStudy (anonymous):

its actually x = 4 got the answer sheet but i dont get how to solve it

OpenStudy (mkmkasim):

nvm it should equal 4...sorry

OpenStudy (mkmkasim):

i added (2/3) you need to subtract

OpenStudy (anonymous):

Lul okay, but can you explain how to solve it?

OpenStudy (anonymous):

please

OpenStudy (mkmkasim):

move 2/3 to the right by subtracting. 6-(2/3)= 16/3

OpenStudy (mkmkasim):

so now you have 4/3x=16/3

OpenStudy (mkmkasim):

no you divide by 4/3, or multiply by 3/4 (same thing) that gives you 16/4=x or more simply put... 4

OpenStudy (anonymous):

oh okay so 2/3 gets cancel out? or what happens to it?

OpenStudy (mkmkasim):

not canceled out... you are subtracting it from both sides...

OpenStudy (anonymous):

ohhhhhh I see

OpenStudy (mkmkasim):